What are the partial pressures And the total pressures of the gases in a 3.7 L balloon, if the balloon is filled with 2.6 moles of gas 1, and 4.8 moles of gas 2 at 58 degrees Celsius?


1
Expert's answer
2021-08-16T02:50:05-0400

number of moles of gas 1 = 2.6 moles

number of moles of gas 2 = 4.8 moles

Total number of moles = 2.6 + 4.8 = 7.4 moles

Temperature = 57°C = (273 + 58)K = 331K


mole fraction of gas 1 = "2.6\/ 7.4 =0.351"

mole fraction of gas 2 = "4.8\/7.4= 0.649"


Using, "PV =nRT"

"P = \dfrac{nRT}{V} = \dfrac{7.4× 8.314× 331}{3.7×10^{-3}}"


"= 5.504MPa"


Partial pressure of 1 "= 0.351\u00d7 5.5MPa" "= 1.93MPa"

Partial pressure of 2 = "0.649 \u00d7 5.5MPa"

"= 3.57MPa"